Knowing that angular frequency is w=SQRT(k/m), what is the predicted angular frequency of -(k1+k2)x=ma?
Answer #1

Here (k1+k2)x = ma can be written in the form of SHM condition with single spring, kx = ma where k = k1+k2.

Now applying the formula for angular frequency w = sqrt(k/m)

= sqrt((k1+k2)/m) answer